OpenStudy (anonymous):

a square has an area of 225 square feet.find the length of each side

OpenStudy (anonymous):

use the square root on a calculator

OpenStudy (anonymous):

15 * 15 = 225

OpenStudy (anonymous):

length of side = sqrt(225) =15 feet

OpenStudy (anonymous):

let the side be x; x ^2=225 so it is x=scrt(225)ft.

OpenStudy (anonymous):